OpenStudy (anonymous):

A survey of 260 families showed that 99 had a dog; 76 had a cat; 34 had a dog and a cat; 98 had neither a cat nor a dog nor a parakeet; 8 had a cat and a dog and a parakeet. How many had a parakeet only? would the answer be 21?

OpenStudy (aum):

Yes.

OpenStudy (aum):



OpenStudy (texaschic101):

nicely done :)

OpenStudy (anonymous):

AUM can you help me with one more question?!?!?!

OpenStudy (aum):

(65 - y) + y + 8 + 26 + (42 - x) + x + z = 216 - 98 141 + z = 162 z = 162 - 141 = 21

OpenStudy (anonymous):

Find the cardinal number of the indicated set by referring to the following table. Find E∩W.

OpenStudy (anonymous):

would it be 234?????

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

aum i didn't pass my test a 5th time do u have a skupe where i can share my screen and u can help me?

OpenStudy (anonymous):

i am really tired.

OpenStudy (aum):

For the previous problem, 234 is the value of the intersection of E and W. Not sure what they mean by cardinal number. Sorry I don't have Skype. The cardinal number may be just 1 as there is just one element, 234, in E intersection W.